# Language-centered Human Activity Recognition

### v3

@@ Abstract.
Human Activity Recognition (HAR) using Inertial Measurement Unit (IMU) sensors is critical for applications in healthcare, safety, and industrial production. However, variations in activity patterns, device types, and sensor placements create distribution gaps across datasets, reducing the performance of HAR models. To address this, we propose LanHAR, a novel system that leverages Large Language Models (LLMs) to generate semantic interpretations of sensor readings and activity labels for cross-dataset HAR. This approach not only mitigates cross-dataset heterogeneity but also enhances the recognition of new activities. LanHAR employs an iterative re-generation method to produce high-quality semantic interpretations with LLMs and a two-stage training framework that bridges the semantic interpretations of sensor readings and activity labels. This ultimately leads to a lightweight sensor encoder suitable for mobile deployment, enabling any sensor reading to be mapped into the semantic interpretation space. Experiments on five public datasets demonstrate that our approach significantly outperforms state-of-the-art methods in both cross-dataset HAR and new activity recognition. The source code will be made publicly available.
@@
copyright: noneconference: ; ; ccs: Human-centered computing Ubiquitous and mobile computing
@@ 1. Introduction
Human Activity Recognition (HAR) based on data collected from Inertial Measurement Unit (IMU) sensors on mobile platforms such as smartphones and wearable devices is one of the key problems in mobile computing due to its important applications in fields such as healthcare (Subasi et al., 2018, 2020), safety (Sun and Chen, 2022; Liagkou et al., 2022), and industrial production (Niemann et al., 2021; Zheng et al., 2018). However, variations in activity patterns, device types, and sensor placements across different individuals result in significant distributional differences between datasets, even when capturing the same activity (Xu et al., 2023; Liu et al., 2022). Consequently, the performance of HAR models deteriorates considerably in cross-dataset human activity recognition scenarios (Xu et al., 2023). Therefore, developing HAR models capable of generalizing across different datasets remains a critical research problem that needs solutions.
@@
We argue that semantic interpretations of sensor readings and activity labels can play a crucial role in addressing the aforementioned limitations as shown in Fig. 1. (i) For cross-dataset heterogeneity, interpreting sensor readings based on its physical meanings can help identify patterns associated with specific activities (e.g., jogging often
@@
exhibits regular periodicity). The underlying intuition is that, despite the variability across datasets, activities may still reveal common underlying patterns. By identifying these patterns, we can map the heterogeneous data into a common space, thereby mitigating the impact of dataset variability. (ii) When dealing with new activities, semantic interpretations of activity labels can enhance their meaning by generating detailed descriptions. Combining semantic interpretations of both sensor readings and activity labels creates the potential to map new data to new activity labels. However, manual semantic interpretation is resource-intensive. The recent success of large language models (LLMs) offers a promising solution, enabling the generation of semantic interpretations from both sensor readings and activity labels through natural language (Jiang et al., 2024; Zhang et al., 2023). Research has shown that LLMs have the ability to perceive aspects of the physical world (Xu et al., 2024), providing an opportunity to leverage these capabilities to bridge the gap in cross-dataset HAR through automated semantic interpretation.
@@
  • To our knowledge, we are the first to leverage LLMs to generate semantic interpretations for cross-dataset human activity recognition. Our method introduces a novel perspective by utilizing semantic interpretations from both sensor readings and activity labels, which helps mitigate the impact of cross-dataset heterogeneity and opens the door to recognizing new activities.
  • We design LanHAR, a novel system that leverages LLMs to generate semantic interpretations of sensor readings and activity labels, addressing the challenge of cross-dataset human activity recognition. The system features a semantic interpretation generation process with an iterative re-generation method to ensure high-quality outputs, alongside a two-stage training framework that transfers the capabilities of large language models to mobile devices.
@@ 2.2. Semantic interpretations of sensor readings for cross-dataset HAR
Activity Raw data Self-supervised learning Semantic interpretations
@@ 2.3. New activity recognition
We argue that combining the semantic interpretations of sensor readings and labels offers a new perspective for addressing this challenge. By converting all activity labels into a language space using their semantic interpretations, we can represent labels even without corresponding sensor readings (which is common for new labels). Similarly, with sensor readings mapped into the same language space, we can match each reading to a label, regardless of whether the label was previously encountered. Fig. 4 illustrates this concept, where we assume “biking” and “going upstairs” are known activities from the training phase, while “going downstairs” is a new activity. In our method, we first convert all activity labels into the same language space using their LLM-generated semantic interpretations. We then obtain semantic interpretations for the sensor readings. Even though “going downstairs” is unseen during training, we can still measure its distance to all activity labels in the language space, enabling the recognition of new activities. In our setting, we assume that we have the knowledge of possible activity labels in the inference stage.
@@ 3.1. Problem formulation
Figure 5. Overview of LanHAR. (1) Utilize LLMs for generating semantic interpretations of sensor reading and activity labels. (2) Train a text encoder to encode two types of semantic interpretations and achieve their alignment. $H{i}$ and $Z_{i}$ denote embeddings of semantic interpretations of activity labels and sensor reading. (3) Train a sensor encoder to align sensor reading and semantic interpretation. $E{i}$ denote embeddings of sensor reading. (4) For inference, only use the sensor encoder to generate embeddings for the sensor readings $E{i}$ and then compute the similarity with the pre-stored embeddings of the activity labels $H{i}$ to obtain the human activity recognition results.
@@ 3.3.1. Obtaining semantics interpretation of sensor reading
(i) Data introduction: We provide details about the data, including its source, the meaning of each dimension, the sampling frequency, and the context in which the data was collected during different activities.
(ii) Data analysis: Intermediate steps (i.e., Chain-of-Thought (Wei et al., 2022)) can enhance LLMs’ reasoning abilities. Our offline testing revealed that while LLMs are capable of performing various data analyses, skipping intermediate steps often leads to less confident responses. To address this, we design multiple auxiliary analysis steps, including amplitude, frequency, time series analysis, and statistical measures like mean, standard deviation, maximum, and minimum values. These steps help LLMs generate more accurate and precise answers.
(iii) Knowledge: We include relevant background knowledge to support the LLMs in making accurate judgments. In many cases, specialized knowledge is essential for accurate interpretation. For example, we describe typical patterns observed in accelerometer and gyroscope readings during activities like running, which aids the LLMs in focusing on the appropriate knowledge area. This knowledge can be sourced from public databases like Wikipedia or generated by LLMs.
@@ 3.3.3. An iterative re-generation method to ensuring the quality of LLM responses
Figure 8. The iterative re-generatio method to ensuring the quality of LLM responses. (1) Filter inaccurate semantic interpretations. (2) Regenerate new semantic interpretations by LLMs. (3) Incorporate the new semantic interpretation based on its impact on the intra-cluster distance.
(1) Filter out inaccurate semantic interpretations: We apply a clustering-based method (i.e., K-means) to identify anomalous data points within the generated semantic interpretations of the same activity. After clustering, we select $k$ outliers based on their distance from the cluster center, with those farther than the average intra-cluster distance considered low-quality responses for refinement.
@@ 3.4.1. Text encoder
We adopt a pre-trained language model as the text encoder due to its robust ability to understand semantic interpretations through language. Specifically, we use the text encoder to encode the semantic interpretations of sensor readings $\mathbf{S}^{z}$ and activity labels $\mathbf{S}^{l}$, which are represented as $\mathbf{Z}=g_{\text{en}}(\mathbf{S}^{z})$ and $\mathbf{H}=g_{\text{en}}(\mathbf{S}^{l})$, where $g_{\text{en}}(\cdot)$ denotes the text encoder. In our work, we utilize the BERT architecture as the text encoder, initializing it with pre-trained BERT parameters (Devlin et al., 2018). BERT consists of multiple Transformer encoder layers, each incorporating multi-head self-attention mechanisms and feedforward neural networks, allowing it to generate word representations in a bidirectional context. We chose BERT because of its superior ability to understand text semantics and compare text similarity, as it captures word meanings in a bidirectional context, enabling nuanced comprehension and effective comparisons. However, this text encoder structure is flexible and can be substituted with other language models if needed.
@@ 3.4.2. Alignment task
$s(\mathbf{H},\mathbf{Z})=\frac{\mathbf{H}}{||\mathbf{H}||}\cdot\frac{\mathbf{Z}}{||\mathbf{Z}||}$
@@ 3.5.1. Sensor encoder structure
The objective of the sensor encoder is to capture key features and generate embeddings from IMU data, which can be formulated as: $\mathit{\mathbf{E}}=f_{sen}(\mathbf{X})$ where $\mathbf{X}\in\mathbb{R}^{S_{\text{dim}}\times L}$ represents the IMU sensor readings, and $\mathbf{E}$ denotes the corresponding embeddings. The structure of the sensor encoder is flexible, allowing for either a CNN or Transformer-based architecture. In our work, we adopt a Transformer-based sensor encoder. First, following (Xu et al., 2021), we normalize the IMU sensor readings $\mathbf{X}$ to obtain the normalized readings, $\mathbf{X^{nor}}$. Next, we project the normalized sensor readings through a linear layer and apply layer normalization to the projected data: $\mathbf{X^{{}^{\prime}}}=\text{LayerNorm}(W_{p}\mathbf{X^{nor}})$ where $W_{p}$ is the weight matrix of the linear layer. Then, we add positional encoding to the normalized data:$\mathbf{X^{0}}=\mathbf{X^{{}^{\prime}}}+\text{PositionEncoding}(\mathbf{X^{{}^{\prime}}})$ The position-encoded data $\mathbf{X^{0}}$ is then fed through the Transformer encoder, which processes the data across multiple layers to capture complex relationships and dependencies. Each Transformer layer consists of two main sub-layers: Multi-Head Self-Attention and a Feed-Forward Network, both followed by layer normalization and residual connections. For the $i$-th encoder layer, the operations are as follows: First, the Multi-Head Self-Attention mechanism calculates self-attention scores, producing the output:
@@
Similar to the alignment task in Section 3.4.2, we adopt contrastive learning to align embeddings of sensor reading with the embeddings of semantic interpretation of sensor reading. First, the embeddings of both the sensor reading and its semantic interpretation are normalized to unit length. Their similarity is calculated using the dot product between the normalized embeddings, denoted as $s(\mathbf{E},\mathbf{Z})=\frac{\mathbf{E}}{||\mathbf{E}||}\cdot\frac{\mathbf{Z}}{||\mathbf{Z}||}.$ The goal of the training is to maximize the similarity between matching pairs while minimizing the similarity between non-matching pairs. Given a batch of N sensor reading and semantic interpretation of sensor reading pairs, the loss function is computed as:

### v4

# Large Language Model-Guided Semantic Alignment for Human Activity Recognition

@@ Abstract.
Human Activity Recognition (HAR) using Inertial Measurement Unit (IMU) sensors is critical for applications in healthcare, safety, and industrial production. However, variations in activity patterns, device types, and sensor placements create distribution gaps across datasets, reducing the performance of HAR models. To address this, we propose LanHAR, a novel system that leverages Large Language Models (LLMs) to generate semantic interpretations of sensor readings and activity labels for cross-dataset HAR. This approach not only mitigates cross-dataset heterogeneity but also enhances the recognition of new activities. LanHAR employs an iterative re-generation method to produce high-quality semantic interpretations with LLMs and a two-stage training framework that bridges the semantic interpretations of sensor readings and activity labels. This ultimately leads to a lightweight sensor encoder suitable for mobile deployment, enabling any sensor reading to be mapped into the semantic interpretation space. Experiments on five public datasets demonstrate that our approach significantly outperforms state-of-the-art methods in both cross-dataset HAR and new activity recognition. The source code is publicly available at https://github.com/DASHLab/LanHAR.
@@
copyright: ccjournal: IMWUTjournalyear: 2025journalvolume: 9journalnumber: 4article: 230publicationmonth: 12doi: 10.1145/3770652ccs: Human-centered computing Ubiquitous and mobile computing
@@ 1. Introduction
Human Activity Recognition (HAR) based on data collected from Inertial Measurement Unit (IMU) sensors on mobile platforms such as smartphones and wearable devices is one of the key problems in mobile computing due to its important applications in fields such as healthcare (Subasi et al., 2018, 2020), safety (Sun and Chen, 2022; Liagkou et al., 2022; Yan et al., 2023), and industrial production (Niemann et al., 2021; Zheng et al., 2018; Yan et al., 2024). However, variations in activity patterns, device types, and sensor placements across different individuals result in significant distributional differences between datasets, even when capturing the same activity (Xu et al., 2023; Liu et al., 2022). Consequently, the performance of HAR models deteriorates considerably in cross-dataset human activity recognition scenarios (Xu et al., 2023). Therefore, developing HAR models capable of generalizing across different datasets remains a critical research problem that needs solutions.
@@
We argue that semantic interpretations of sensor readings and activity labels can play a crucial role in addressing the aforementioned limitations as shown in Fig. 1. (i) For cross-dataset heterogeneity, interpreting sensor readings based on its physical meanings can help identify patterns associated with specific activities (e.g., jogging often exhibits regular periodicity). The underlying intuition is that, despite the variability across datasets, activities may still reveal common underlying patterns. By identifying these patterns, we can map the heterogeneous data into a common space, thereby mitigating the impact of dataset variability. (ii) When dealing with new activities, semantic interpretations of activity labels can enhance their meaning by generating detailed descriptions. Combining semantic interpretations of both sensor readings and activity labels creates the potential to map new data to new activity labels. However, manual semantic interpretation is resource-intensive. The recent success of large language models (LLMs) offers a promising solution, enabling the generation of semantic interpretations from both sensor readings and activity labels through natural language (Jiang et al., 2024; Zhang et al., 2023). Research has shown that LLMs have the ability to perceive aspects of the physical world (Xu et al., 2024), providing an opportunity to leverage these capabilities to bridge the gap in cross-dataset HAR through automated semantic interpretation.
@@
  • We explore the use of LLM-generated semantic representations to improve cross-dataset human activity recognition. Our method maps both sensor readings and activity labels into a shared semantic space using natural language descriptions. This alignment reduces dataset heterogeneity and enables the recognition of activities that are not present in the training data.
  • We design LanHAR, a novel system that leverages LLMs to generate semantic interpretations of sensor readings and activity labels, addressing the challenge of cross-dataset human activity recognition. The system features a semantic interpretation generation process with an iterative re-generation method to ensure high-quality outputs, alongside a two-stage training framework that transfers the capabilities of large language models into lightweight, privacy-preserving models suitable for sustainable deployment on resource-constrained edge devices.
@@
2. Related work
2.1. Human activity recognition
Human Activity Recognition (HAR) leverages data from various sensors such as accelerometers and gyroscopes to detect and classify human activities (Yao et al., 2017; Yang et al., 2015). To mitigate the issue of data heterogeneity, existing works can be classified into three categories, including semi-supervised learning methods, domain adaptation methods, and data augmentation methods.
In HAR, collecting large-scale labeled datasets is particularly challenging due to the high cost of manual annotation, while abundant unlabeled sensor data are readily available. To address this imbalance, semi-supervised methods (Xu et al., 2021; Saeed et al., 2019; Zhang et al., 2024; Miao and Chen, 2024; Presotto et al., 2023) have been developed to maximize the use of unlabeled data by learning valuable feature representations without requiring explicit labels. These approaches enable models to adapt more efficiently to downstream tasks with only a minimal amount of labeled data, which is crucial for improving generalization across diverse datasets. For example, UniMTS (Zhang et al., 2024) is a unified pre-training framework designed for motion time-series data, with zero-shot and few-shot generalization across diverse human activity recognition tasks. LIMU-BERT (Xu et al., 2021) leverages masked language modeling and self-supervised learning to effectively utilize unlabeled IMU sensor data. Despite their capabilities, these models still depend on a certain amount of labeled data to train classifiers, which may struggle when applied to target datasets that lack any labeled data. GOAT (Miao and Chen, 2024) uses pre-training with textual attributes to improve activity recognition across multiple datasets, supporting cross-dataset generalization.
The core idea of domain adaptation methods (Chang et al., 2020; Qin et al., 2019; Khan et al., 2018; Zhou et al., 2020) involves taking a classifier that has been pre-trained on a source domain and using it with an unlabeled dataset from a target domain, then adjusting the source model’s weights to enhance its performance in the target domain. For example, SDMix (Lu et al., 2022) aims to mix features from the source and target domains in the feature space while emphasizing the preservation of semantic discrimination, in order to reduce the discrepancy between different domains. UDAHAR (Chang et al., 2020) systematically studies the effectiveness of various unsupervised domain adaptation methods for human activity recognition tasks. It benchmarks multiple UDA techniques, such as feature alignment and adversarial adaptation, providing a comprehensive evaluation of their robustness under different domain shifts. While domain adaptation methods mitigate domain shifts, they often struggle with large domain discrepancies, depend heavily on abundant source data
Data augmentation (Um et al., 2017; Xu et al., 2023; Qian et al., 2022; Saeed et al., 2019; Wang et al., 2022; Tang et al., 2021) in human activity recognition aims to increase the diversity and quality of training data, thereby enhancing the model’s generalization ability and accuracy. For example, UniHAR (Xu et al., 2023) designs physical-informed augmentation methods to improve the cross-dataset HAR performance.
However, All the methods are valuable, they still suffer from two limitations: (i) Most existing work overlooks the physical meanings of activities. This can exacerbate the impact of classification errors on downstream applications. For instance, misclassifying a walking activity as running may have a relatively minor effect compared to misclassifying it as a significantly different activity such as biking. (ii) Moreover, cross-dataset scenarios often introduce new activities that these methods struggle to handle, as these new activities are typically missing from the training phase. Our potential idea to solve these limitations is to use and map semantic interpretations of sensor readings and activity labels in a shared semantic space to mitigate cross-dataset heterogeneity and generate new activity.
Recognizing new activities has been a longstanding challenge in HAR, particularly for machine learning-based methods, which are unable to identify activities that were not present during the training phase. The most relevant method (Cheng et al., 2013) attempts to address the challenge of recognizing new activities by using pre-defined attributes, such as actions of specific body parts. The method trains two mappings: one from sensor readings to predefined attributes, and another from predefined attributes to activity labels. However, the quality of these predefined attributes significantly influences the model performance. Additionally, this work is not well-suited for cross-dataset new activities, as the new activities may differ significantly from the original dataset.
2.2. LLM for human activity recognition
Recently, there has been a surge of interest in leveraging the capabilities of LLMs (Naveed et al., 2025; Jin et al., 2023; Tan et al., 2025) to enhance mobile computing applications (Lee et al., 2024; Wen et al., 2024). Penetrative AI (Xu et al., 2024) explores the potential of leveraging large language models to comprehend and interact with the physical world. Kim et al. (Kim et al., 2024) harness the power of large language models to improve health prediction using data collected from wearable sensors. IMUGPT2.0 (Leng et al., 2024) utilizes LLMs for virtual IMU data generation since IMU-labeled data is challenging to collect. It incorporate motion filtering and diversity metrics to enhance the efficiency and practicality of synthetic data generation for HAR. MotionGPT (Jiang et al., 2023) fuses language data with large-scale motion models, and motion-language pre-training that can enhance the performance of motion-related tasks becomes feasible. HARGPT (Ji et al., 2024) leverages LLMs as zero-shot human activity recognizers by framing sensor-based activity recognition as a natural language processing task. ContextGPT (Arrotta et al., 2024) uses LLMs to extract commonsense knowledge about the environmental contexts of human activities, enhancing context-aware HAR with limited supervision.
Recent works exploring the use of large language models for human activity recognition have demonstrated the potential of LLMs to enhance HAR in various ways and have inspired our approach to leverage LLMs for generating semantic interpretations of IMU data. Nevertheless, while many studies explore LLMs in the context of HAR, a few directly address HAR tasks, and even those that still face challenges in enabling efficient deployment on resource-constrained mobile devices.
@@ 3.2. Semantic interpretations of sensor readings for cross-dataset HAR
Activity
Raw data
Self-supervised
learning
Semantic
interpretations
@@ 3.3. New activity recognition
We argue that combining the semantic interpretations of sensor readings and labels offers a new perspective for addressing this challenge. By converting all activity labels into a language space using their semantic interpretations, we can represent labels even without corresponding sensor readings (which is common for new labels). Similarly, with sensor readings mapped into the same language space, we can match each reading to a label, regardless of whether the label was previously encountered. Fig. 4 illustrates this concept, where we assume “biking” and “going upstairs” are known activities from the training phase, while “going downstairs” is a new activity. In our method, we first convert all activity labels into the same language space using their LLM-generated semantic interpretations. We then obtain semantic interpretations for the sensor readings. Even though “going downstairs” is unseen during training, we can still measure its distance to all activity labels in the language space, enabling the recognition of new activities. In our setting, we assume that we have the knowledge of possible activity labels in the inference stage. It is important to note that our approach does not consider open-set recognition or arbitrary label discovery. Instead, we follow a closed-set assumption, where the candidate activity labels are known in advance during inference. To support this, we pre-stored all semantic interpretations of activity labels and use them to guide activity recognition during inference. Note that this does not violate our “new activity” objective where we do not have access to the training data of new activities, instead, simply the “name” of the activities.
@@ 4.1. Problem formulation
Figure 5. Overview of LanHAR. (1) Utilize LLMs for generating semantic interpretations of sensor reading and activity labels. (2) Train a text encoder to encode two types of semantic interpretations and achieve their alignment. HiH{i} and $Z_{i}$ denote embeddings of semantic interpretations of activity labels and sensor reading. (3) Train a sensor encoder to align sensor reading and semantic interpretation. EiE{i} denote embeddings of sensor reading. (4) For inference, only use the sensor encoder to generate embeddings for the sensor readings EiE{i} and then compute the similarity with the pre-stored embeddings of the activity labels HiH{i} to obtain the human activity recognition results.
@@ 4.3.1. Obtaining semantics interpretation of sensor reading
(i) Data introduction: We provide details about the data, including its source, the meaning of each dimension, the sampling frequency, and the context in which the data is collected during different activities.
(ii) Data analysis: Intermediate steps (i.e., Chain-of-Thought (Wei et al., 2022)) can enhance LLMs’ reasoning abilities. Our offline testing reveals that while LLMs are capable of performing various data analyses, skipping intermediate steps often leads to less confident responses. To address this, we design multiple auxiliary analysis steps (e.g., amplitude, frequency, time series analysis, and statistical measures such as mean, standard deviation, maximum, and minimum values). These steps help LLMs generate more accurate and precise answers.
(iii) Knowledge: We include relevant background knowledge to support the LLMs understanding the analysis procedure and focusing on the critical aspects among many possible features (e.g., impact peaks and vertical–horizontal dominance). The background knowledge encompasses a few key aspects of sensor pattern characteristics and highlights their importance in distinguishing between different specific activities or broader activity categories. The background knowledge is mainly derived from established findings in prior research on IMU-based human activity recognition (Kavanagh and Menz, 2008; Nurwulan and Selamaj, 2021; Chen et al., 2021; Wang et al., 2016; Barna et al., 2019; Li et al., 2009; Wang et al., 2011). For example, we referred to studies on how vertical and horizontal components reflect different movement mechanisms, how inertial signals vary across activities, and how dominant gyroscope axes can indicate trunk rotation or arm-swing patterns. When needed, we also consulted open sources (e.g., Wikipedia) for basic terminology. All information was then consolidated through a combination of LLM-based summarization and manual refinement.
@@ 4.3.3. An iterative re-generation method to ensuring the quality of LLM responses
Figure 8. The iterative re-generation method to ensuring the quality of LLM responses. (1) Filter inaccurate semantic interpretations. (2) Regenerate new semantic interpretations by LLMs. (3) Incorporate the new semantic interpretation based on its impact on the intra-cluster distance.
(1) Filter out inaccurate semantic interpretations: We first encode the generated semantic interpretations into dense vector representations using a pre-trained BERT encoder. We then apply a clustering-based method (i.e., K-means) to identify anomalous data points within the generated semantic interpretations of the same activity. After clustering, we select $k$ outliers based on their distance from the cluster center, with those farther than the average intra-cluster distance considered low-quality responses for refinement.
@@ 4.4.1. Text encoder
We adopt a pre-trained language model as the text encoder due to its robust ability to understand semantic interpretations through language. Specifically, we use the text encoder to encode the semantic interpretations of sensor readings $\mathbf{S}^{z}$ and activity labels $\mathbf{S}^{l}$, which are represented as $\mathbf{Z}=g_{\text{en}}(\mathbf{S}^{z})$ and $\mathbf{H}=g_{\text{en}}(\mathbf{S}^{l})$, where $g_{\text{en}}(\cdot)$ denotes the text encoder. In our work, we utilize the BERT architecture as the text encoder, initializing it with pre-trained BERT parameters (Devlin et al., 2018). BERT consists of multiple Transformer encoder layers, each incorporating multi-head self-attention mechanisms and feedforward neural networks, allowing it to generate word representations in a bidirectional context. We chose BERT because of its superior ability to understand text semantics and compare text similarity, as it captures word meanings in a bidirectional context, enabling nuanced comprehension and effective comparisons. Unlike generative models that process text in a left-to-right (or autoregressive) manner, BERT considers both the left and right context of each word simultaneously during training, allowing it to build deeper semantic representations based on the full sentence structure. However, this text encoder structure is flexible and can be substituted with other language models if needed.
@@ 4.4.2. Alignment task
[…]
@@ 4.5.1. Sensor encoder structure
[…]
@@
[…]